OpenStudy (anonymous):

test the convergence or divergence, sum (n=0..inf)(1/2n-1/2n+6)

OpenStudy (anonymous):

help me answer this questions

OpenStudy (amoodarya):

\[\sum_{1}^{\infty}\frac{ 1 }{ 2n }-\frac{ 1 }{ 2n+6 }=\frac{ 1 }{ 2 }\sum_{1}^{\infty}\frac{ 1 }{ n }-\frac{ 1 }{ n+3 }=\\\frac{ 1 }{ 2 }\sum_{1}^{\infty}\frac{ 1 }{ n }-\frac{ 1 }{ n+1 }+\frac{ 1 }{ n+1 }-\frac{ 1 }{ n+2 }+\frac{ 1 }{ n+2 }-\frac{ 1 }{ n+3 }=\]

OpenStudy (amoodarya):

now you have 3 sum in form of f(n)-f(n+1) so now put n=1,2,3,... simplify you will see it converge

OpenStudy (anonymous):

ouw....thank you...

OpenStudy (amoodarya):

understand it ?

OpenStudy (anonymous):

yes....

OpenStudy (anonymous):

the equation is telescoping series right?